OpenStudy (anonymous):

Graph the following linear inequality: y <= 7 x -1

OpenStudy (anonymous):

The plots I got are: y=-1 & x=-1/7?

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

So Close! :D 0 = 7x - 1 +1 +1 1 = 7x /7 /7 x = 1/7

OpenStudy (anonymous):

I did it that way too, but i don't know why I choose the negative one lol

OpenStudy (anonymous):

How do I graph fractions? :/

jigglypuff314 (jigglypuff314):

hmmmm... there's another way to get an easier point to graph... you know how in y = mx + b m = slope? so in this case slope = 7 which means that for every time you move to the right one, it goes up 7

OpenStudy (anonymous):

so is 7 the rise or run?

jigglypuff314 (jigglypuff314):

slope = rise over run = 7 over 1 so 7 is the rise :)

OpenStudy (anonymous):

ok, so I go from y=-1, then go 7 up and 1 over?

jigglypuff314 (jigglypuff314):

yes :) you have points (0, -1) and (1, 6)

OpenStudy (anonymous):

Thank you, next question.